Help I don't know how to solve this.
-6=x/15

Answer by TimothyLamb(4379)

-6 = x / 15
x / 15 = -6
15(x / 15) = 15*(-6)
(15/15)x = 15*(-6)
x = -90
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php